### v1

# Traditional Methods Outperform Generative LLMs at Forecasting Credit Ratings

Abstract.
Large Language Models (LLMs) have been shown to perform well for many downstream tasks. Transfer learning can enable LLMs to acquire skills that were not targeted during pre-training. In financial contexts, LLMs can sometimes beat well-established benchmarks. This paper investigates how well LLMs perform in the task of forecasting corporate credit ratings. We show that while LLMs are very good at encoding textual information, traditional methods are still very competitive when it comes to encoding numeric and multimodal data. For our task, current LLMs perform worse than a more traditional XGBoost architecture that combines fundamental and macroeconomic data with high-density text-based embedding features. The code from this paper is provided [^1] https://github.com/FelixDrinkall/credit-ratings-project.
Forecasting, NLP, LLM, Credit Ratings, Generative, XGBoost, Clustering, Multimodal
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; pre-print; July 10th, 2024
1. Introduction
Corporate credit ratings indicate a borrower’s ability to service its debt obligations and are a forward-looking measure of a company’s health (Baresa et al., 2012). A company’s credit rating is significant since it affects the cost of raising capital, which in turn could finance future infrastructure to increase revenue or profitability. An optimistic rating can result in a virtuous cycle whereby it is easier to raise money and grow the business (Cho et al., 2020), and a pessimistic rating can result in a vicious cycle in which competition can grow faster due to cheaper debt obligations. Knowing which cycle a company may enter can be advantageous to investors. Many major funds are also not allowed to own sub-prime assets, which makes forecasting a drop in credit rating very important so that the fund has more time to divest from the asset, which could result in a higher close price.
Figure 1. Example of the best-performing feature-type - high-density clustering (Drinkall et al., 2022). Each dot represents a sentence, and the colored dots represent sentences that exist in high-density regions of embedding space.
Recently, there has been a surge of interest in text-based forecasting (Xu and Cohen, 2018; Yang et al., 2020; Nie et al., 2024). One reason for this trend is the progress that has been made in text modelling in general (Wei et al., 2022; Touvron et al., 2023). Given that financial news is often first disseminated through written or spoken communications (Boulland et al., 2016), rather than in numeric or tabular formats, there has been a hope that important information can be included in models sooner than was possible without using linguistic information. Another reason is that language can provide relevant context and forward-looking information, whereas financial numeric reporting alone is inherently retrospective. Contained within a company’s filings, text can provide insights about the future strategic direction of the company as well as historical information.
The majority of text-based forecasting research has been directed at analyzing short-sequence textual content, drawing primarily from sources such as social media (Xu and Cohen, 2018), news articles (Zhang et al., 2018), and analyst recommendations (Rekabsaz et al., 2017). In contrast, many fundamental financial documents, such as company filings, earnings call transcripts, and patents, are very long. Considering that the shorter texts often serve as summaries, reflections or commentaries on the detailed primary sources and that speed of information acquisition is essential in finance (Rzayev et al., 2023), there must be more focus on text-based forecasting for longer text sequences. This paper evaluates the most effective ways to model longer text sequences within a text-based forecasting task.
Linked to the recent progress of LLMs, there has been growing interest in applying language models in a wide variety of downstream applications (Kaddour et al., 2023). It has been shown that as generative LLMs scale, they acquire abilities that were not present in smaller LLM variants (Wei et al., 2022), such as modular arithmetic (Srivastava et al., 2023), NLU (Hendrycks et al., 2021), commonsense reasoning (Lin et al., 2022), fact-checking (Rae et al., 2022) and so on. These abilities have been impressive, but it is best not to be over-optimistic. The lack of training data transparency associated with some of the best-performing language models means that we cannot be certain whether some of the performance gains are due to the memorisation of benchmarks being in the training datasets (Bender et al., 2021; Sainz et al., 2023; Xu et al., 2024; Balloccu et al., 2024). Generative language models also seem to have a mediocre understanding of concepts like negation and complex logical reasoning (Kassner and Schütze, 2020; Lorge and Pierrehumbert, 2023; Huang and Chang, 2023; Truong et al., 2023). These limitations in the capabilities of language models could prove to be very consequential in financial contexts. In this paper, we test generative language models on a complex linguistic task, which has never been fully solved by human experts: credit rating forecasting. We show that while language models encode text-based information very well, they are not good at incorporating numeric information, and underperform a boosting-tree baseline.
1.1. Contributions
  • We show that generative LLMs are poor at encoding numerical information, and underperform traditional methods.
  • To our knowledge, this is the first use of modern language modelling techniques in a credit rating forecasting task.
  • A financial dataset that can be reproduced with an academic WRDS licence.
  • A benchmark of different techniques for encoding long-sequence text in a forecasting task.
2. Related Work
2.1. Text-based forecasting
2.1.1. Encoding Text for Forecasting
The predominant approach in text-based forecasting has focused on the extraction of interpretable features like sentiment and uncertainty scores (Song and Shin, 2019; An et al., 2023). Rule-based sentiment using diverse lexicons has dominated the literature (Mohammad, 2020; Kalamara et al., 2022; Barbaglia et al., 2023). Lexicons tailored to specific domains generally surpass broader lexicons in predictive tasks (Loughran and McDonald, 2011; Li et al., 2014). Nevertheless, lexicons overlook contextual nuance and inadequately address common linguistic phenomena like negation. To mitigate these limitations, efforts have been made to integrate more sophisticated sentiment classifiers (An et al., 2023; Ayyappa et al., 2023). However, sentiment presupposes that important information can be encapsulated within a single dimension. To avoid an overly simple and prescriptive feature set, unsupervised methods have been used in feature exploration: TF-IDF (Jones, 1972), Latent Dirichlet Allocation (LDA) (Wang et al., 2017; Kanungsukkasem and Leelanupab, 2019). However, the arrival of contemporary topic models has gradually eclipsed LDA, fostering the adoption of transformer-derived topic models into forecasting tasks (Drinkall et al., 2022).
Recently, some studies have used the representations from encoder-based LLMs as features for text-based forecasting. LLMs exploit high-dimensional embeddings to capture the linguistic meaning of words (Devlin et al., 2019; Radford et al., 2019) and sentences (Reimers and Gurevych, 2019), with representation dimensionality ranging from 384 (Wang et al., 2020) to 5192 (Touvron et al., 2023). Such dimensionality poses challenges when used with smaller datasets. Nonetheless, there has been some success in incorporating these methods into text-based forecasting (Sawhney et al., 2020; Lee et al., 2023). However, the effectiveness of LLMs is often hampered by their limited context windows. Recent advancements have seen an increase in context window sizes, thanks in part to better GPU infrastructure making it computationally feasible, the implementation of attention sparsification techniques (Tay et al., 2022), and positional encoding hacks (Chen et al., 2023). There are also methods that combine the use of transformer-based LLMs with feature-based methods, such as topic clusters (Grootendorst, 2022; Drinkall et al., 2022), or emotions (Liapis and Kotsiantis, 2023).
2.1.2. Generative Multimodal Forecasting
In addition to the adoption of encoder-based LLMs like BERT (Devlin et al., 2019), generative LLMs have been used in text-based forecasting tasks. Generative LLMs use masked-self attention to model text in an autoregressive manner. The GPT (Radford et al., 2019) and Llama (Touvron et al., 2023) model families are part of the generative model class. LLMs have been used as a backbone model for generative time-series forecasting models (Cao et al., 2023; Chang et al., 2023; Zhou et al., 2023; Liu et al., 2024), showing that an adapted generative language model can forecast weather, electricity and several other domains without relying on traditional text inputs. (Liu et al., 2024) used eight text-based frames in order to create a general time-series modal that could be applied to several domains, and in so doing encoded both text and numerical information in a GPT-2-small model to generate the predicted time-series.
Beyond the use of text-based frames in generative forecasting tasks, GPT4MTS (Lee et al., 2023) encoded both news and time-series information before passing the concatenated input sequence through a pre-trained GPT-2 model. FinMA (Xie et al., 2023) and PromptCast (Xue and Salim, 2023) evaluated the performance of language models on stock movement prediction by converting the time-series information into natural language and prompting the language model for the predicted direction. (Yu et al., 2023) takes this further by passing exclusively text information into the prompt for a financial forecasting task. There has been little comparison between these generative methods and the more traditional discriminative methods when applied to multimodal information.
2.2. Credit Rating Prediction
Research in Credit Rating Prediction (CRP) has tended to focus on predicting the absolute credit rating at time $t=0$ given the feature set $F_{t=0}$ (Li et al., 2023; Galil et al., 2023; Tavakoli et al., 2023). This approach takes the perspective of the rating agencies and is useful for identifying anomalies where the existing credit rating classification appears to be implausible or inconsistent with current financial indicators (Lokanan et al., 2019). However, predicting the absolute rating level is more simple and not as useful as predicting a future change. There is some limited research on Credit Rating Forecasting (CRF), where the target is the movement direction of the credit rating at time $t=1$. This task takes the perspective of the investor seeking to predict whether an asset is likely to be classified as more or less risky in the next time period, and is the task outlined in this paper.
There have been some attempts to incorporate linguistic information into both corporate risk (Fei et al., 2015; Cao et al., 2024) and default prediction (Mai et al., 2019; Stevenson et al., 2021). Some papers have shown how text can help improve consumer credit lending (Hurley and Adebayo, 2016; Babaei and Giudici, 2024). There has also been some attempts to include textual data in CRP and CRF tasks (Chen and Chen, 2022; Muñoz-Izquierdo et al., 2022; Tavakoli et al., 2023). The majority of the existing literature uses lexicons, keywords or sentiment to encode the text (Kogan et al., 2009; Fei et al., 2015; Mai et al., 2019; Muñoz-Izquierdo et al., 2022; Chen and Chen, 2022). There have been some studies that have utilised encoder-based LLM representations (Stevenson et al., 2021; Tavakoli et al., 2023; Cao et al., 2024). There has been some work exploring how well generative models perform at assessing credit lending applications (Babaei and Giudici, 2024), and value at risk in general (Cao et al., 2024), but there has been no work benchmarking how well modern generative language models perform on a CRF task. Understanding generative LLMs’ relative strengths relative to more traditional methods is an important contribution to the existing literature.
3. Dataset
In part due to the lack of large open-source or readily available datasets with temporal metadata, most of the financial text-based forecasting studies have either focused on expensive proprietary datasets, or datasets spanning 2-3 years (Xu and Cohen, 2018; Soun et al., 2022), making results hard to replicate and potentially biased to a specific time. While temporal bias in language-based tasks is hard to avoid due to limited historical data (Drinkall et al., 2024), we aim to reduce this by using a dataset spanning 23 years, enhancing the generalizability of our findings across different economic contexts. The cost and lack of transparency of large datasets have hindered progress in the field and made it harder to build on promising work due to the difficulty of replicating results.
3.1. Data Sources
All data used is either open source or available with a WRDS subscription to enable effective dataset reconstruction. The data used in this paper is derived from exclusively US-based companies.
3.1.1. Credit ratings (CR)
For the credit ratings we used the Compustat Capital IQ dataset [^2] Credit Ratings: https://tinyurl.com/r4urtkc5, using Standard & Poors’ (S&P) ratings. These ratings cover the period from 1978 to 2017. S&P routinely assesses and assigns credit ratings to companies. Our paper predicts changes to the long-term credit ratings. Notably, we incorporate historical ratings from preceding quarters into our prediction models, acknowledging the distinct implications of a top-rated company (AAA) being downgraded compared to a lower-rated one (CC) experiencing a similar decline.
3.1.2. SEC filings
This paper utilizes 10-Q and 10-K filings available in the SEC’s EDGAR database [^3] Filings Database: https://www.sec.gov/edgar/searchedgar/companysearch to provide both textual context. They were chosen for their consistent structure which aids homogenous feature extraction. While most of the content in these filings is comprised of indexing, tables, and introductory text, we’re interested in the parts that offer insights into a company’s future financial health. As such, we’ve focused on the Management’s Discussion and Analysis of Financial Condition and Results of Operations (MDA) section. This paper extracts the MDA sections from all SEC filings - using the SEC-API [^4] Filings API: https://sec-api.io/ - for which we had credit rating data, spanning from Q1 of 1994 to Q2 of 2017. The API itself returns cleaned text, but we clean the text further by removing the remaining HTML, links and excessive spaces.
3.1.3. Fundamental data (F)
S&P emphasize two primary components in their credit rating methodology: the financial and business risk profiles (Gillmor, 2015). Whilst the text from the MDA section can provide some insight into the qualitative business risk profile, numerical fundamental data is important to assess the financial health of a company. For the fundamental data, we use the Compustat Quarterly Fundamentals dataset [^5] Fundamental Data:https://tinyurl.com/4ca8ddst. The variables selected are outlined in Appendix D. Importantly, these variables were consistently reported for all the companies under consideration. Ideally, we would incorporate a broader range of fundamental variables, but expanding the variable set would result in fewer samples with complete data, thus limiting the scope of our analysis.
3.1.4. Macroeconomic data (M)
Adverse events in the world economy can also impact a company’s ability to repay its debt. Many external forces can affect a company’s future creditworthiness, however, we have identified three key areas from prior research in the area (Carling et al., 2007; Taylor et al., 2021): labour statistics, interest rates and foreign exchange data. For the labour statistics, we used the Bureau of Labour Statistics dataset [^6] Labour Statistics: https://tinyurl.com/y94d52xk. For the interest rate and foreign exchange data, we used the Federal Reserve Bank Reports [^7] Interest Rate Data: ${}^{\text{,}}$ [^8] Foreign Exchange Data: https://tinyurl.com/a38rmzd8.
3.2. Dataset Construction
To ensure consistency with the periodicity of SEC filings, all data in our study is aligned every quarter. This dataset extends from Q1 1994, marking the beginning of the SEC’s electronic processing of filings, to Q2 2017, the last point at which we have credit rating data from Compustat Capital IQ. Not every company in our dataset has a complete and continuous record of data. Consequently, we have excluded instances with incomplete data. A significant outcome of this approach is that when the number of lagged quarters used in the task is increased, the number of valid samples diminishes. This reduction is due to the lower probability of having complete data across many consecutive quarters, compared to when only the most recent quarter is considered.
The class imbalance in credit rating data is highly skewed. 93.4% of the companies in our dataset maintain the same score. Whilst it is common convention to use oversampling techniques such as SMOTE (Chawla et al., 2002) to aid credit rating prediction models (Pamuk and Schumann, 2023; Wang, 2022; Zhao et al., 2024). There is no consensus on how SMOTE could be applied to high-dimensional text embeddings. As a result, we balanced the classes for this task, reducing the size of the datasets. The train test splits run from 01-01-1994 to 31-12-2012, the validation splits run from 01-01-2013 to 31-12-2014, and the test splits from 01-01-2015 to 31-12-2016.
The MDA section of an SEC Filing, despite only constituting a small part of the filing, is still very long. The average MDA section considered in our task is 13,267 tokens long using a BPE tokenizer (Sennrich et al., 2016). As a result, when a model could not encode all of the tokens in the MDA section, only the first part of the text was encoded, up to either a set length or the maximum input sequence length.
4. Methodology
We deploy two frameworks to test different architectural methodologies on this task We maintain the same data available to each variation of the frameworks. The first framework is a feature-based discriminative approach that uses a more traditional boosting-tree model and tests the different ways to encode the textual data. The second uses generative language models and prompting to output one of a fixed list of labels through a greedy search algorithm (Appendix B).
4.1. Task Description
The objective of the task is to predict the credit rating, $\hat{R}_{t}$, at time $t$. The function can be represented as follows:
$\hat{R}_{t}=f\big{(}$ $T_{t-1},T_{t-2},\ldots,T_{t-p};$
$R_{t-1},R_{t-2},\ldots,R_{t-p};$
$N_{t-1},N_{t-2},\ldots,N_{t-p}\big{)}$
Here, $T_{t-i}$ represents the text data, $R_{t-i}$ represents the historical credit rating data, $N_{t-i}$ represents the numeric data - both fundamental and macroeconomic. $i$ varies from 1 to $p$, with $p$ indicating the number of past quarters considered (1 to 4 quarters in this study). Furthermore, $f$ is the predictive function to convert our input data into an estimate. An ablation study is conducted to evaluate the impact of different data types on the prediction accuracy. In this study, the function $f$ is tested under various configurations: using only text data $T_{t-i}$, using combinations of historical ratings $R_{t-j}$, and numeric data $N_{t-k}$. This approach helps to determine the relative importance of each type of data.
4.2. Boosting-Tree Baseline
To test the abilities of generative language models, it is necessary to benchmark the performance against a relatively well-understood and robust algorithm. We select XGBoost (Chen and Guestrin, 2016), a model that has been widely adopted in many domains (Talukder et al., 2023; Dong et al., 2023; Joshi et al., 2024). The supervised model takes as input the normalised fundamental, macroeconomic and text data, and outputs the most likely label.
4.3. Text Encoders
Data type Av. Quarters
1 2 3 4
Numeric M + F + CR 52.81 48.29 53.32 53.99 55.65
CR 44.68 41.92 43.57 46.48 46.73
All LM Lex 50.64 46.83 51.19 52.13 52.42
Emotion Lex 51.35 49.76 52.41 51.60 51.61
LDA 50.88 50.25 52.34 50.76 50.16
Clusters 53.61 50.73 54.64 54.13 55.96
BERT Emotion 52.83 48.23 52.92 55.35 54.83
BERT 52.69 50.51 54.09 52.29 53.89
Text only LM Lex 34.42 36.56 33.42 30.85 36.83
Emotion Lex 35.51 32.93 34.48 36.17 38.44
LDA 34.75 36.59 35.01 30.32 37.10
Clusters 38.12 39.76 38.01 38.86 35.83
BERT Emotion 36.01 36.83 35.81 36.17 35.22
BERT 32.65 32.20 28.91 36.17 33.33
[…]
4.3.1. Loughran McDonald Lexicon (LM Lex)
(Loughran and McDonald, 2011) - The LM Lexicon is widely recognized in finance. Given its prevalence, it is crucial to compare its effectiveness with more advanced methods. The lexicon classifies words into four domains: Positivity, Negativity, Litigiousness, and Uncertainty. However, the simple language modelling technique classifies phrases like ”The debt increased last quarter” as neutral. The LM text representation in this work is the count of how many words from each sentiment occur in the document, normalized by the maximum value in the train set.
4.3.2. NRC Lexicon (Emotion Lex)
(Mohammad, 2020) - The NRC lexicon is a general lexicon which scores words for ten different emotions and sentiments and is used for multi-domain problems. The text representation is created in the same way as the LM Lexicon.
4.3.3. Latent Dirichlet Allocation (LDA)
(Blei et al., 2003) A prominent topic modelling technique often employed to assign topics to sequences of text. Although the field of topic modelling has evolved, LDA continues to serve as a robust baseline for evaluating new topic models. To create the text representation we created a vector where each of the dimensions represented the probability that the text belonged to each one of 25 different topics.
4.3.4. High-density Embedding Clusters (Clusters)
(Grootendorst, 2022) - Contextual word embeddings have provided a good basis for topic modelling (Sia et al., 2020; Grootendorst, 2022). Sentence embeddings have also been used to discern domain type from text (Aharoni and Goldberg, 2020). (Drinkall et al., 2022) extended this work to generate features from clusters of sentence embeddings in a COVID-19 caseload prediction task. For this task, each sentence of each filing in the train set was encoded into embeddings space using a all-mpnet-base-v2 (Reimers and Gurevych, 2019), the dimensionality was then reduced using UMAP (McInnes et al., 2020), and the HDBSCAN clustering algorithm (Campello et al., 2013) was used for form 100 distinct clusters. Examples of the clusters will be included in supplementary material. Then each filing in the train, validation and test set was split into sentences and then transformed into the embedding space described above. The overall text representation was the average of the representations of each sentence, and the representation of each sentence was the probability distribution that the sentence belonged to each of the 100 clusters.
4.3.5. DistilRoBERTa Emotion (BERT Emotion)
(Hartmann, 2022) - In this work a DistilRoBERTa model has been fine-tuned on an emotion classification task that includes Ekman’s basic emotions (Ekman, 1992), plus the neutral label: anger, disgust, fear, joy, neutral, sadness, surprise. This methodology provides a more complex comparison of whether emotions are a useful feature base for predicting a change in credit ratings. The SEC Filings are then chunked into 512 token sequences and classified according to the probability that that chunk can be associated with each emotion. The average across all chunks is taken as the final text representation of each filing.
4.3.6. BERT
(Devlin et al., 2019) - The first 512 tokens of each SEC Filing are passed into BERT and the average embedding from the final layer is taken as the textual representation of that filing.
Model Data type Average Quarters Feature Importances
1 2 3 4 Macro Fund Text
XGBoost Numeric 52.81 48.29 53.32 53.99 55.65 0.6593 0.3407 -
All (Clusters) 53.86 50.73 54.64 54.13 55.96 0.1174 0.1880 0.6946
Text Only (Clusters) 38.12 39.76 38.01 38.86 35.83 - - -
Numeric Features + GPT4o Text Estimate 53.81 52.68 50.93 53.99 57.26 0.5877 0.3023 0.1010
XGBoost Numeric Estimate + GPT4o Text Estimate 51.75 53.66 51.19 50.80 51.34 0.5727 0.4273
GPT-4o Numeric 31.41 33.66 31.57 31.12 29.30 - - -
All 40.24 43.90 40.32 38.83 37.90 - - -
Text Only 49.64 49.27 52.23 52.39 44.62 - - -
Text + XGBoost Numeric Estimate 32.27 33.90 30.77 36.70 27.69 - - -
Union of XGBoost-Numeric & GPT-4o Text-Only 69.87 70.49 73.21 71.54 64.25 - - -
[…]
4.4. Generative Framework
Generative LLMs have seen significant recent progress. Given this development, it is important to see if these models can discern whether there is likely to be a change in the perceived risk of a company, and if so, what is the best methodology for achieving high performance. This section is methodologically distinct from the other text encoding methods described above since the numerical data is encoded into text before the model attempts to solve the task through the use of prompting. The prompts used in the following section are included in the Appendix F, and follow the best practice from existing literature (OpenAI, 2024; Lin et al., 2024).
Whilst LLMs perform very well in 0-shot settings (Kojima et al., 2023), there is significant evidence that shows that LLMs perform better in a k-shot setting (Clark et al., 2018); the ARC benchmark uses 25-shot prompts in the Eleuther AI evaluation harness (Gao et al., 2023). The problem with deploying a k-shot framework in this setting is that the SEC Filings are very long (13,267 tokens). Despite the increase in the context-window length of some newer LLMs, many new models are capped at 8192 tokens or below (Wang and Komatsuzaki, 2021; Jiang et al., 2023; Touvron et al., 2023; Singh et al., 2024), and some other studies have shown performance deterioration as the input sequence increases (Li et al., 2024). Fitting several examples of the task in the input sequence is impossible for many of the data samples, which means that k-shot performance is not reported for this task.
We tested the following models using the prompting structure laid out in Appendix F: gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09 and gpt-4o-2024-05-13, Llama-3 8B. These models provide a good representation of the current state-of-the-art (Chiang et al., 2024).
4.4.1. LoRA Adaptation
To adapt the LLMs to the task we use LoRA (Low Rank Adaptation) (Hu et al., 2021) fine-tuning. This technique involves optimizing the rank-decomposition matrices, $A$ $B$, of the change in model weights ($\Delta W$), where $W^{{}^{\prime}}$ are the new model weights and $W$ are the pre-trained weights.
(1) $W^{{}^{\prime}}$ $=W+\Delta W$
(2) $=W+BA$
The advantage is that it requires a lot less memory to fine-tune a model and in contrast to some parameter-efficient fine-tuning methods, adaptation can take place through the entire model stack.
Data Model Av. Quarters
1 2 3 4
Num. Llama-3 8B 32.31 35.12 35.81 28.99 29.30
Llama-3 8B LoRA 35.51 35.37 34.22 37.77 34.68
GPT-3.5 32.63 32.93 33.42 31.91 32.26
GPT-4 34.08 34.15 32.89 32.98 36.29
GPT-4o 31.41 33.66 31.57 31.12 29.30
All Llama-3 8B 35.50 35.37 35.81 36.97 33.87
Llama-3 8B LoRA 37.47 36.39 37.01 37.64 38.83
GPT-3.5 44.52 49.02 42.44 46.01 40.59
GPT-4 38.28 39.76 36.07 38.03 39.25
GPT-4o 40.24 43.90 40.32 38.83 37.90
Text Llama-3 8B 35.57 35.37 36.07 36.97 33.87
Llama-3 8B LoRA 36.96 38.10 36.75 37.11 35.87
GPT-3.5 46.36 47.32 47.48 45.48 45.16
GPT-4 48.48 47.80 48.54 50.27 48.12
GPT-4o 49.64 49.27 52.23 52.39 44.62
[…]
5. Results
The results from the XGBoost baseline are outlined in Table 1. Firstly, it is clear that there is some information in the text since almost all text encoding methods perform above chance. None of the individual text feature sets outperform the numeric baselines, showing that the fundamental and macroeconomic variables are more important to the prediction than the text. However, when all of the features are combined we see a performance increase when the high-density cluster features are used in conjunction with the numerical information.
The generative models show some interesting behaviour in Table 3. Using a 0-shot prompt, the performance using numerical information is near random. Indeed the performance is higher on the GPT-class models when only text is used as opposed to when all data types are used. This shows that the numeric information is harming the predictive performance. Interestingly, GPT-3.5 - an older model - performs the best when all data is considered. It also appears that LoRA enables slightly better relative performance on understanding numerical data - Llama-3 8B LoRA is the best-performing model on entirely numerical information and is the only model with no performance degradation when all features are considered as opposed to just text. However, the main conclusion from Table 3 is that generative models are very good relatively at decoding text data for this task.
Table 2 takes the best-performing text features from the XGBoost framework, Clusters, and compares the model to the best-performing generative model, GPT-4o. Interestingly, GPT-4o utilises the text alone much better than any of the other encoder-based methods, however when all features are considered the XGBoost-Clusters configuration is the best-performing methodology. In addition to that, the models are also picking up on different signals. The proportion of samples where at least one of the XGBoost-Numeric and GPT-4o Text Only is correct (69.87) is significantly higher than any of the individual models. As a result, we provide comparisons where the estimate and probability of the XGBoost-Numeric and GPT-4o Text-Only are included in the prompt or feature set. Both the estimate and the probability that each model assigns to the estimate are used in the prompt or as features. None of the estimate combination methods outperform the XGBoost-Clusters framework, but it seems reasonable to suggest that future ensemble methods may be able to capture some of this performance gap.
6. Interpretability
One of the major disadvantages of using generative LLMs is that, to a large degree, they are black boxes. While some work has successfully used attention weights to analyse generated prompts (Serrano and Smith, 2019; Moradi et al., 2019), many of the best-performing LLMs are only accessible via APIs (Chiang et al., 2024), which remove the ability to analyse LLMs’ attention weights. Boosting trees, using known and interpretable features, on the other hand, provide increased interpretability. Regulation in major economies is moving towards a system that emphasizes explainability almost as much as absolute performance (Commission, 2020; US Congress, 2022; UK Secretary of State for Digital, Culture, Media and Sport, 2022); methods that provide a window into the statistical associations that the model infers during its decision-making process are essential. As well as providing the best performance of all the model configurations in this paper, the XGBoost-Clusters framework provides end-users with the ability to understand and explain the decisions that made. This section provides an example of how we can use this framework to understand the reasons behind decisions.
(a) PDP of ”rating_sp_ratings_moodys” cluster & ”Up” class.
(b) PDP of ”receivables_percent_owned_offs” cluster & ”Down” class.
Figure 2. Partial Dependence Plots (PDP) of text-based features against different target classes.
The most obvious advantage of a feature-based system is that important features can be identified. Table 2 provides an example of feature importance that can be used to infer the modality preference of model configurations. It is possible to conduct even more granular feature analysis by looking at the contribution of individual features. Figure 2 shows the partial dependence plot of some of the individual text features on the ”Up” and ”Down” classes. From the plot in Figure 2(a) we can infer that as ratings are discussed more in a company’s filing, there is a reduced chance of the the credit rating being upgraded. We can also infer from Figure 2(b) that as companies talk about receivables - the money owed to the company - there is a reduced chance of the credit rating being downgraded. Both provide valuable insights, and are examples of how a traditional feature-based methodology can be leveraged.
7. Conclusion
The paper shows that while LLMs are good at encoding textual data and inferring signals that traditional methods cannot pick up on, when combined with numerical information in the prompt there is performance deterioration. Traditional methods still outperform LLMs at combining multimodal data. The other advantage is that traditional methods offer increased interpretability and a better understanding of the mechanisms behind certain predictions.
There has been some work jointly encoding text using generative LLMs with time-series information (Liu et al., 2024), but more work needs to be done to determine the best methodology for combining long text sequences with numerical information while utilizing the benefits of generative LLM natural language understanding. This paper shows that combining multimodal information within the prompt is not sufficient.
Appendix A Limitations
The task above uses a balanced dataset, which is good for testing the different methodologies’ ability to discern the signals that are predictive of a rise or fall in credit ratings, but is poor for assessing how good the models would be in a real-world context where almost all of the ratings stay the same. As such, despite the data being taken from across all US equities over a 23-year time-period, the balanced dataset is relatively small, with only 3441 samples for the Lag 1 configuration and 2142 samples for the Lag 4 configuration. There are plenty of prominent datasets that are smaller, but the size reduces the scope for complex and specialised models to be deployed on this task in favour of more robust, simple models.
Another limitation is that the text used in this paper is produced by the company, who have an agenda to convey a positive agenda to investors. More objective publication venues may produce different insights about the future direction of a company.
We also assume that the credit rating methodology remains the same between the train and test sets. This is an assumption that is made by the rest of the literature, and our training set is spread over an 18 year period, however it does not rule out the possibility that the results are only valid over the time period that was tested. Due to the size of the dataset we were restricted from using a masked temporal cross-validation evaluation framework, which would have left insufficient data for training for some years.
The LoRA fine-tuning methodology outlined in this paper is a parameter-efficient technique and has been shown to be competitive in a variety of settings (Hu et al., 2021), but can be outperformed in some tasks by full fine-tuning and other adapter-based methods (Xu et al., 2023). We compared the performance of the LoRA implementation in this paper to that of QLoRA (Dettmers et al., 2023), which produced marginally worse results. However, it is possible that other fine-tuning techniques would have produced better results. Further to this, it is possible that if a better model than Llama-3 8B had been fine-tuned we would have seen even better results from the generative LLMs. The computational constraints placed on us are not dissimilar to those that other researchers face, which makes the results in this paper valid while perhaps not exhaustive.
Appendix B Greedy Decoding
Generative models can produce unpredictable outputs (Fadaee and Monz, 2020; Stureborg et al., 2024), which necessitates the use of constrained generation when an LLM forms part of a larger architecture. For the purposes of this work, we use a greedy search to infer the probability of one of the following labels appearing next in the sequence: ”up”, ”down” or ”same”.
Appendix C Dataset Size
The size of each of the dataset splits are outlined in Table 4.
# Quarters Train Val Test
1 2,642 389 410
2 1,748 374 377
3 1,595 351 376
4 1,445 325 372
Table 4. Dataset sizes
Appendix D Fundamental Data
The fundamental variables considered are outlined in Table 5.
Variable Type Description
niq Float Net Income (Loss)
ltq Float Liabilities - Total
piq Float Pretax Income
atq Float Assets - Total
ggroup Char GIC Groups
gind Char GIC Industries
gsector Char GIC Sectors
gsubind Char GIC Sub-Industries
Table 5. Description of Variables
Appendix E Neural Network Implementations
We also tested some more complex neural network (NN) approaches, which had underwhelming results. The Hierarchical Credit Rating (HierCR) model is a framework that models the filings hierarchically. The challenge with using language models to encode the filings is the limited context window of encoder-based language models. There have been many different solutions to this problem, including sparse attention mechanisms (Beltagy et al., 2020), chunking (Sawhney et al., 2020), and feature-based extraction like the methods above (Loughran and McDonald, 2011; Drinkall et al., 2022). Our NN solution to this problem is to split the filing up into sentences and pass the sentence embeddings through an all-mpnet-base-v2 encoder to produce embeddings for the textual data. The text encoder replicates the structure in (Sawhney et al., 2020), the only material difference is that filing sentences substitute the social media posts in the first layer of the text-encoder. The text, macro and fundamental vectors across the previous quarter(s) are combined using a GRU layer (Chung et al., 2014), the outputs are then passed through an attention layer to create a representation for each data-type. These representations are combined using a bilinear transformation, which is passed through 3 linear layers followed by a ReLU (Agarap, 2019) activation function. Dropout is applied in the final linear-layer classification module.
Model Av. Quarter
1 2 3 4
SA-LF 40.62 39.87 40.05 41.17 41.39
SA-EF 43.41 41.24 40.97 45.31 46.11
HierCR 35.02 33.58 32.89 35.62 37.98
Table 6. Accuracy for more complex NN approaches using all data types (M+F+T).
The other two architectures are Shared Attention Late Fusion (SA-LF) and Shared Attention Early Fusion (SA-EF). Both architectures only consider the first 512 tokens of each filing. The difference between the two architectures is when the attention layer is applied. For the SA-EF the model attends to all feature-types together, whereas the SA-LF only combines the representations after the attention layer is applied to the individual data-types. The final representation is the passed through the same linear-layer classification module as the HierCR. For all of the architectures above, we trained the model for 200 epochs with a patience value of 20 epochs.
The results from these models are poor in comparison to the more simple XGBoost models. This could be due to the size of the dataset, which does not provide the model enough data to train on without overfitting the training data. Complex models with many parameters require more data to fit properly. Given that the dataset is the largest balanced and complete dataset possible to make using US data, and that the size of the dataset considered in this paper is representative of a large number of other tasks, the results from this paper represent a significant contribution for dealing with problems of this nature.
Appendix F Prompts
To encode the numerical and textual information into text form, we used the prompting structure outlined below. When the ablation study was carried out the prompt and data included was adjusted accordingly.
### System: You are trying to work out whether a company’s credit rating is likely to go up, down, or stay the same given its recent credit ratings. Predict the likely movement in a company’s credit rating for the next quarter, using historical credit ratings, quantitative financial data and macroeconomic data. The numeric data has been normalized and appears in order with the most recent first.
### Credit Rating Explanation:
Credit ratings use the following scale, in order of increasing risk: ’AAA’, ’AA+’, ’AA’, ’AA-’, ’A+’, ’A’, ’A-’, ’BBB+’, ’BBB’, ’BBB-’, ’BB+’, ’BB’, ’BB-’, ’B+’, ’B’, ’B-’, ’CCC’, ’CCC-’, ’CC’, ’C’, ’SD’
### Fundamental Financial Indicators Defined:
### Macroeconomic Variables Defined:
### User:
Your task is to classify the company into one of the following classes: ”down”, ”same”, ”up”. ”down” means that you think the credit rating will go down in the next quarter, meaning the company is perceived as more risky. ”same” means that you think the credit rating will stay the same in the next quarter. ”up” means that you think the credit rating will go up in the next quarter, meaning the company is perceived as less risky. Please respond with a single label that you think fits the company best.
Classify the following numerical data:”””
F.1. Credit Rating Ranking
One potential problem with the prompt outlined in Appendix F is that the LLM may find it hard to correctly understand the ranking structure of credit ratings, which would limit the ability of an LLM to perform well on this task. To probe the LLMs ability to understand the relative rank of credit ratings we created the following prompt:
”””Two credit ratings will be given, the task is to determine which is higher on the following scale, which is ordered in descending order:
’AAA’, ’AA+’, ’AA’, ’AA-’, ’A+’, ’A’, ’A-’, ’BBB+’, ’BBB’, ’BBB-’, ’BB+’, ’BB’, ’BB-’, ’B+’, ’B’, ’B-’, ’CCC’, ’CCC-’, ’CC’, ’C’, ’SD’.
Please answer with the higher rating e.g. AAA vs. SD Answer: AAA.
¡¡rating_X¿¿ vs. ¡¡rating_Y¿¿ Answer:”””
The performance on this task across all rating combinations when prompting GPT-4o was 99.52%. The only mistake was between C and CC. This high performance displays a very good understanding of the credit rating scale and justifies the setup of our prompt.
Appendix G S&P Credit Rating Definitions
S&P’s definitions for each of the credit rating categories are outlined in Table 7.
Category Definition
AAA An obligation rated ’AAA’ has the highest rating assigned by S&P Global Ratings. The obligor’s capacity to meet its financial commitment on the obligation is extremely strong.
AA An obligation rated ’AA’ differs from the highest-rated obligations only to a small degree. The obligor’s capacity to meet its financial commitment on the obligation is very strong.
A An obligation rated ’A’ is somewhat more susceptible to the adverse effects of changes in circumstances and economic conditions than obligations in higher-rated categories. However, the obligor’s capacity to meet its financial commitment on the obligation is still strong.
BBB An obligation rated ’BBB’ exhibits adequate protection parameters. However, adverse conditions or changing circumstances are likely to lead to a weakened capacity of the obligor to meet its financial commitment on the obligation.
BB; B; CCC; CC; and C Obligations rated ’BB’, ’B’, ’CCC’, ’CC’, and ’C’ are regarded as having significant speculative characteristics. ’BB’ indicates the least degree of speculation and ’C’ the highest. While such obligations will likely have some quality and protective characteristics, these may be outweighed by large uncertainties or major exposures to adverse conditions.
BB An obligation rated ’BB’ is less vulnerable to nonpayment than other speculative issues. However, it faces major uncertainties or exposure to adverse business, financial, or economic conditions which could lead to the obligor’s inadequate capacity to meet its financial commitment on the obligation.
B An obligation rated ’B’ is more vulnerable to nonpayment than obligations rated ’BB’, but the obligor currently has the capacity to meet its financial commitment on the obligation. Adverse business, financial, or economic conditions will likely impair the obligor’s capacity or willingness to meet its financial commitment on the obligation.
CCC An obligation rated ’CCC’ is currently vulnerable to nonpayment, and is dependent upon favorable business, financial, and economic conditions for the obligor to meet its financial commitment on the obligation. In the event of adverse business, financial, or economic conditions, the obligor is not likely to have the capacity to meet its financial commitment on the obligation.
CC An obligation rated ’CC’ is currently highly vulnerable to nonpayment. The ’CC’ rating is used when a default has not yet occurred, but S&P Global Ratings expects default to be a virtual certainty, regardless of the anticipated time to default.
C An obligation rated ’C’ is currently highly vulnerable to nonpayment, and the obligation is expected to have lower relative seniority or lower ultimate recovery compared to obligations that are rated higher.
SD An obligation rated ’SD’ is in default or in breach of an imputed promise. For non-hybrid capital instruments, the ’SD’ rating category is used when payments on an obligation are not made on the date due, unless S&P Global Ratings believes that such payments will be made within five business days in the absence of a stated grace period or within the earlier of the stated grace period or 30 calendar days.
NR This indicates that no rating has been requested, or that there is insufficient information on which to base a rating, or that S&P Global Ratings does not rate a particular obligation as a matter of policy.
Table 7. S&P Global Ratings Definitions (spg, 2016)